# DeepMIDE: A Multivariate Spatio-Temporal Method for Ultra-Scale Offshore Wind Energy Forecasting

### v1

@@ Abstract
To unlock access to stronger winds, the offshore wind industry is advancing with significantly larger and taller wind turbines. This massive upscaling motivates a departure from univariate wind forecasting methods that traditionally focused on a single representative height. To fill this gap, we propose DeepMIDE—a statistical deep learning method which jointly models the offshore wind speeds across space, time, and height. DeepMIDE is formulated as a multi-output integro-difference equation model with a multivariate, nonstationary, and state-dependent kernel characterized by a set of advection vectors that encode the physics of wind field formation and propagation. Embedded within DeepMIDE, an advanced deep learning architecture learns these advection vectors from high-dimensional streams of exogenous weather information, which, along with other parameters, are plugged back into the statistical model for probabilistic multi-height space-time forecasting. Tested on real-world data from future offshore wind energy sites in the Northeastern United States, the wind speed and power forecasts from DeepMIDE are shown to outperform those from prevalent time series, spatio-temporal, and deep learning methods.
@@ 1 Introduction
Offshore wind energy is poised to play a major role in the worldwide transition towards cleaner and more sustainable energy systems (Musial et al., 2023). To harness stronger winds, the offshore wind industry is advancing with the next-generation offshore wind turbines, which are going to be significantly taller and larger than conventional wind turbines. With hub heights exceeding $140$ meters (m) and rotor diameters as large as $220$m, these ultra-scale turbines are set to become the largest rotating machines on earth, approaching the altitude of many of the world’s tallest landmarks (Gaertner et al., 2020; General Electric, 2021; Zahle et al., 2024). By producing significantly higher per-capita energy output using the same wind resources, considerable reductions in the cost of harnessing wind energy resources can be realized relative to conventional wind turbine designs that typically have had hub heights and rotor diameters of $100$m or less (U.S. Department of Energy, 2021; Shields et al., 2021).
@@
While the massive upscaling of offshore wind turbines unlocks significant benefits related to the “economy of scale,” it also raises considerable challenges pertaining to the optimal operation of those massive-scale assets in a highly dynamic and uncertain environment. Today, the reliable management of offshore wind farms heavily relies on short-term wind speed and power forecasts to mitigate the uncertainty of the intermittent wind resources (Giebel et al., 2016). Those forecasts have been traditionally issued using statistical models that describe the temporal evolution of wind conditions at a single representative height, potentially combining information from multiple geographical locations (Sweeney et al., 2020). By doing so, those models compress the three-dimensional reality of the wind field (space $\times$ time $\times$ height) into a simpler, two-dimensional “world view” (space $\times$ time) which completely overlooks the vertical dimension of the wind field. In this paper, “space” refers specifically to the horizontal dimensions (latitude and longitude), while “height” refers to the vertical dimension (altitude).
@@
While this simplification may be tenable for conventional wind turbines, it is indeed not plausible for the next-generation of offshore wind turbines—there is growing evidence in the wind power engineering literature that larger-scale wind turbines experience a significant influence from vertical wind shear on the turbine’s power production. Hence, the power output of those ultra-scale generators more accurately correlates with the full vertical wind profile swept by the now-much-larger rotor diameter, rather than with a single horizontal slice representing the average hub-height wind speeds (Antoniou et al., 2009; Wagner et al., 2011; Barthelmie et al., 2020). This observation motivates us to seek more sophisticated models for offshore wind forecasting in which the evolution of wind conditions at multiple locations and heights are described in tandem. Our survey of the literature, presented in Section 2, reveals that there is very limited work on multi-height space-time wind forecasting. We believe that, until recently, this has been largely due to the lack of a practical need for such models. Evidently, this is no longer the case with the emergence of ultra-scale wind turbines.
@@
As we argue in this paper, making the leap from single- to multi-height space-time wind forecasting is not a trivial task, primarily due to the complex and non-separable dependencies of the full-fledged wind field across the space-time-height trio. To fill this gap, we propose a multi-output spatio-temporal model with a multivariate, nonstationary, and state-dependent kernel which—unlike existing space-time wind forecasting approaches—is capable of adequately modeling the complex dependencies in local wind speeds across multiple heights. Embedded within the statistical model, an advanced deep learning architecture fully exploits the wealth of exogenous information available to the forecaster (typically in the form of high-dimensional image streams) in order to learn and update key physically meaningful kernel parameters that encode physically relevant information about wind field dynamics. In that way, our proposed approach retains the desirable properties of a statistical model, while embedding a deep learning architecture to act as an internalized “physics extractor” that exploits the large amounts of high-dimensional exogenous information that is difficult to process using standard statistical machinery. Experiments using real-world data from the offshore wind energy areas in the Northeastern United States show that the wind speed and power forecasts from the proposed model are of considerably higher quality than those from prevalent time series, deep learning, and space-time methods.
@@
The remainder of this paper is organized as follows. Section 2 reviews the literature on space-time wind forecasting in the context of wind energy applications. Section 3 presents the real-world data used in this study and its relevance to the rising offshore wind sector in the Northeastern United States. In Section 4, we present the building blocks of the proposed model, followed by Section 5 where forecast evaluations, results, and findings are discussed. Section 6 concludes the paper and highlights future research directions.
@@ 2 Literature review
The engineering literature on wind turbine modeling acknowledges the influence of vertical wind speed variations on the performance of large-scale wind turbines (Antoniou et al., 2009; Wagner et al., 2011; Lee et al., 2015; Barthelmie et al., 2020; International Electrotechnical Commission, 2022). This literature primarily focuses on modeling a turbine’s power output, conditional on a fully known wind speed profile.
@@
In practice, however, wind speed—the key determinant of wind power output—is the major source of uncertainty in wind energy harnessing, and is not known to wind farm operators in advance. Thus, to optimally operate an (offshore) wind farm, accurate wind speed and power forecasts are needed at various forecast horizons, ranging from a few minutes up to several days ahead. Those forecasts directly inform critical decisions made by energy producers and grid operators alike, including power production estimation (Zhu et al., 2021; Nasery and Ezzat, 2023), electricity market participation and trading (Pinson, 2013), operations and maintenance scheduling (Papadopoulos et al., 2021, 2024b), economic dispatch and unit commitment (Xie et al., 2013; Barry et al., 2022), among others. Despite the general consensus in the wind engineering community on the importance of considering the vertical wind profile, there are very few studies in the forecasting literature on multi-height wind forecasting (Lin and Liu, 2020; Saxena et al., 2021). Even these studies are only focused on a single location, thereby ignoring the spatial variations and correlations in the wind field.
@@
On the other hand, there is an extensive body of literature on space-time wind forecasting, which focuses on modeling the spatial and temporal variations and correlations, but ignores the vertical dimension of wind fields. Perhaps the most prevalent class of models are geostatistical approaches, with spatio-temporal Gaussian processes (GPs) at their heart. GPs make space-time forecasts that are enabled by a covariance function describing the dependence structure over space and time (Gneiting, 2002; Stein, 2005; Ezzat, 2020). A central challenge in GP-based models is to propose a mathematically permissible covariance function that adequately models the complex dependencies in wind fields. One particulary important aspect of that complexity is to model the effect of time-varying wind advection, which gives rise to an asymmetry in the spatial-temporal correlations. This complexity requires a covariance model that relaxes the convenient, (but unrealistic) assumptions of stationarity, separability, and full space-time symmetry (Gneiting et al., 2006a; Ezzat et al., 2018, 2019).
@@
An alternative paradigm to geostatistical approaches are the so-called dynamic spatio-temporal models (DSTMs) which describe the conditional evolution of a spatial process over time (Wikle and Hooten, 2010). The most prevalent example is the integro-difference equation (IDE) framework, which models the dependence between a spatial process at a future time conditioned on the same process at the present time through an integral operator with a spatial redistribution kernel (Wikle and Cressie, 1999). IDE models have been proposed to predict spatio-temporal cloud movement (Wikle, 2002), precipitation (Ke Xu and Fox, 2005; Liu et al., 2022), aerosols (Calder et al., 2011), and more recently, wind power (Ye and Ezzat, 2024). There are pros and cons for the geostatistical and dynamic space-time approaches, with the former typically regarded as superior surrogate models, whereas the latter are well-suited for extrapolation of highly dynamic processes due to their explicit conditional evolution mechanisms.
@@
Different in principle than the first two statistical approaches (i.e., geostatistical models and DSTMs), deep learning methods have demonstrated significant promise in wind forecasting applications by developing architectures that are tailored to extract complex patterns from spatio-temporal wind data (Wang et al., 2021), including models based on convolutional neural networks (CNNs) (Zhu et al., 2019), recurrent neural networks (RNNs) (Ghaderi et al., 2017), graph neural networks (GNNs) (Khodayar and Wang, 2019; Li et al., 2023), and more recently, transformer models (Bentsen et al., 2023). Despite their demonstrated predictive performance, deep learning still falls short in terms of uncertainty quantification and interpretability. Thus, recent lines of research seek hybrid methods, referred to collectively as statistical deep learning approaches, in which elements or characteristics of a deep learning architecture are embedded within a statistical model to perform certain learning tasks. For example, deep Kalman filters (Krishnan et al., 2015) and deep state-space models (Rangapuram et al., 2018) learn the parameters of a state-space model using multi-layer perceptrons (MLPs), and recurrent neural networks (RNNs), respectively. Deep autoregressive models (DeepAR) entail a sequence of RNNs that are structured in an autoregressive-inspired manner (Salinas et al., 2020), wheras deep IDE models invoke a CNN to learn the IDE redistribution kernel parameters using past process observations (De Bézenac et al., 2019; Zammit-Mangion and Wikle, 2020).
@@
Departing from the existing literature, we propose a multi-output IDE model characterized by a multivariate, nonstationary, and state-dependent redistribution kernel that is tailored to capture the complex dependencies and interactions of local wind fields across the space-time-height trio. To exploit the wealth of exogenous information available to the forecaster (typically in the form of high-dimensional images), we embed an advanced deep learning architecture within the multivariate IDE model in order to learn key physically meaningful kernel parameters that encode physically relevant information about local wind field dynamics. The proposed model is therefore statistical in nature, enabling full probabilistic inference and forecasting across space, time, and height, with the internalized deep learning architecture acting as a “physics extractor,” to learn a low-dimensional set of physically relevant parameters using high-dimensional data that are difficult to process using standard statistical machinery.
@@
Our proposed model departs from the vast majority of space-time wind forecasting models (be it geostatistical, DSTMs, or deep-learning-based) that adopt a two-dimensional world view of the wind field (space $\times$ time) by focusing on a single representative height. It belongs to the class of statistical deep learning models reviewed above, and especially builds on the deep IDE models therein (De Bézenac et al., 2019; Zammit-Mangion and Wikle, 2020), but differs in two key aspects. First, it generalizes the univariate IDE approaches to the multivariate setting to model the complex and time-varying dependencies across space, time, and height. In that way, the proposed model is regarded as the multivariate extension of the univariate IDE model, akin to how multi-output GPs generalize their univariate counterparts. Second, unlike existing deep IDE models which use information that is endogenous to the process (e.g., past observations) to train simple deep learning architectures (typically a CNN), our model exploits the wealth of exogenous information available to the forecaster, typically in the form of high-dimensional image streams. For example, when forecasting wind speeds, one has access to large amounts of high-dimensional data from regional numerical weather models about the meso-scale wind velocities, pressure and temperature systems, etc. that encode rich physical information about the local wind dynamics. To fully unlock access to this exogenous information, we propose to embed a transformer model—an advanced deep learning architecture with evidenced success in computer vision and high-dimensional data mining (Vaswani et al., 2017)—within the multivariate IDE framework as a “physics extractor,” i.e., to learn and update a low-dimensional set of time- and height-varying parameters encoding the physics of local wind field formation and propagation.
@@ 3 Data description and analysis
This work is motivated by the ongoing large-scale offshore wind developments in the Northeastern United States, and in particular, the NY/NJ Bight (Bureau of Ocean Energy Management, 2021). Several of those offshore wind projects have already announced plans to install ultra-scale offshore wind turbines. We make use of two sources of data with varying spatial and temporal resolutions: (i) A set of wind speed measurements collected, at multiple heights, by three floating lidar buoys; and (ii) a set of numerical weather predictions (NWPs), of multiple meteorological variables, obtained from a regional numerical weather model. Details of both sets of data are described next.
Figure 1: (a) Locations of the three sites, E05N, E06, and ASOW6, on top of the offshore wind energy areas in the NY/NJ Bight. The background map is generated using the Northeast Ocean Data Portal (Northeast Ocean Data Portal, 2023); (b) Wind rose plot for the wind speeds, averaged over space and height; (c) Time series of wind speeds at all locations (different colors) and heights (different markers), showing strong spatial, temporal, and vertical dependencies, and motivating the need for a joint modeling approach across space, time, and height, as the one advocated in this paper.

### v2

# DeepMIDE: A Multi-Output Spatio-Temporal Method for Ultra-Scale Offshore Wind Energy Forecasting

@@ Abstract
To unlock access to stronger winds, the offshore wind industry is advancing towards significantly larger and taller wind turbines. This massive upscaling motivates a departure from wind forecasting methods that traditionally focused on a single representative height. To fill this gap, we propose DeepMIDE—a statistical deep learning method which jointly models the offshore wind speeds across space, time, and height. DeepMIDE is formulated as a multi-output integro-difference equation model with a multivariate nonstationary kernel characterized by a set of advection vectors that encode the physics of wind field formation and propagation. Embedded within DeepMIDE, an advanced deep learning architecture learns these advection vectors from high-dimensional streams of exogenous weather information, which, along with other parameters, are plugged back into the statistical model for probabilistic multi-height space-time forecasting. Tested on real-world data from offshore wind energy areas in the Northeastern United States, the wind speed and power forecasts from DeepMIDE are shown to outperform those from prevalent time series, spatio-temporal, and deep learning methods.
@@ 1 Introduction
Offshore wind energy is poised to play a major role in the worldwide transition towards cleaner and more sustainable energy systems (musial2023offshore). To harness stronger winds, the offshore wind industry is advancing with the next-generation offshore wind turbines, which are going to be significantly taller and larger than conventional wind turbines. With hub heights exceeding $140$ meters (m) and rotor diameters as large as 220220m, these ultra-scale turbines are set to become the largest rotating machines on earth, approaching the altitude of many of the world’s tallest landmarks (gaertner2020definition; GE; 22mwturbine). By producing significantly higher per-capita energy output using the same wind resources, considerable reductions in the cost of harnessing wind energy can be realized relative to conventional wind turbine designs that typically have had hub heights and rotor diameters of 100100m or less (doefuture; shields2021impacts).
@@
While the massive upscaling of offshore wind turbines unlocks significant benefits related to the “economy of scale,” it also raises considerable challenges pertaining to the optimal operation of those massive-scale assets in a highly dynamic and uncertain environment. Today, the reliable management of offshore wind farms heavily relies on short-term wind speed and power forecasts to mitigate the uncertainty of the intermittent wind resources (giebel2016wind). Those forecasts have been traditionally issued using statistical models that describe the temporal evolution of wind conditions at a single representative height, potentially combining information from multiple geographical locations (sweeney2020future). By doing so, those models compress the three-dimensional reality of the wind field (space $\times$ time $\times$ height) into a simpler, two-dimensional “world view” (space $\times$ time) which completely overlooks the vertical dimension of the wind field. In the context of this paper, “space” refers specifically to the horizontal dimensions (latitude and longitude), while “height” refers to the vertical dimension (altitude).
@@
While this simplification may be tenable for conventional wind turbines, it is indeed not plausible for the next-generation of offshore wind turbines—there is growing evidence in the wind power engineering literature that larger-scale wind turbines experience a significant influence from vertical wind shear on the turbine’s power production. Hence, the power output of those ultra-scale generators more accurately correlates with the full vertical wind profile swept by the now-much-larger rotor diameter, rather than with a single horizontal slice representing the average hub-height wind speeds (antoniou2009wind; wagner2011accounting; barthelmie2020power). This observation motivates us to seek more sophisticated models for offshore wind forecasting in which the evolution of wind conditions at multiple locations and heights are described in tandem. Our survey of the literature, presented in Section 2, reveals that there is very limited work on multi-height space-time wind forecasting. We believe that, until recently, this has been largely due to the lack of a practical need for such models. Evidently, this is no longer the case with the emergence of ultra-scale wind turbines.
@@
As we argue in this paper, making the leap from single- to multi-height space-time wind forecasting is not a trivial task, primarily due to the complex and non-separable dependencies of the full-fledged wind field across the space-time-height trio. To fill this gap, we propose a multi-output spatio-temporal model with a multivariate, nonstationary kernel, which—unlike existing space-time wind forecasting approaches—is capable of adequately modeling the complex dependencies in local wind speeds across multiple heights. Embedded within the statistical model, an advanced deep learning architecture fully exploits the wealth of exogenous information available to the forecaster (typically in the form of high-dimensional image streams) in order to learn and update key physically meaningful kernel parameters that encode physically relevant information about wind field dynamics. In that way, our proposed approach retains the desirable properties of a statistical model, while embedding a deep learning architecture to act as an internalized “physics extractor” that exploits the large amounts of high-dimensional exogenous information that is difficult to process using standard statistical machinery. Experiments using real-world offshore wind data from the Northeastern United States—a region with significant offshore wind energy potential—demonstrate that the wind speed and power forecasts from the proposed model are of considerably higher quality than those from prevalent time series, deep learning, and space-time methods.
@@
The remainder of this paper is organized as follows. Section 2 reviews the literature on space-time forecasting for wind energy applications. Section 3 presents the real-world data used in this study and its relevance to the planned offshore wind energy developments in the Northeastern United States. In Section 4, we present the building blocks of the proposed model, followed by Section 5 where forecast evaluations, results, and findings are discussed. Section 6 concludes the paper and highlights future research directions.
@@ 2 Literature review
The engineering literature on wind turbine modeling acknowledges the influence of vertical wind speed variations on the performance of large-scale wind turbines (antoniou2009wind; wagner2011accounting; lee2015power; barthelmie2020power; iec2022). This literature primarily focuses on modeling a turbine’s power output, conditional on a fully known wind speed profile.
@@
In practice, however, wind speed—the key determinant of wind power output—is the major source of uncertainty in wind energy harnessing, and is not known to wind farm operators in advance. Thus, to optimally operate an (offshore) wind farm, accurate wind speed and power forecasts are needed at various forecast horizons, ranging from a few minutes up to several days ahead. Those forecasts directly inform critical decisions made by energy producers and grid operators alike, including power production estimation (zhu2021multi; nasery2023yaw), electricity market participation and trading (Pinson2013), operations and maintenance scheduling (petros2021; papadopoulos2023joint), economic dispatch and unit commitment (xie2013short; barry2022risk), among others. Despite the general consensus in the wind engineering community on the importance of considering the vertical wind profile, there are very few studies in the forecasting literature on multi-height wind forecasting (lin2020wind; saxena2021offshore). Even these studies are only focused on a single location, thereby ignoring the spatial variations and correlations in the wind field.
@@
On the other hand, there is an extensive body of literature on space-time wind forecasting, which focuses on modeling the spatial and temporal variations and correlations, but ignores the vertical dimension of wind fields. A relevant class of models therein are geostatistical approaches, such as spatio-temporal Gaussian processes (GPs). GPs make space-time forecasts that are enabled by a covariance function describing the dependence structure over space and time (GneitingCov; SteinCov; AzizGP). A central challenge in GP-based models is to propose a mathematically permissible covariance function that adequately models the complex dependencies in wind fields. An important aspect of that complexity is to model the effect of time-varying wind advection, which gives rise to an asymmetry in the spatial-temporal correlations. This complexity requires a covariance model that relaxes the convenient (but unrealistic) assumptions of stationarity, separability, and full space-time symmetry (gneiting2006geostatistical; AzizRS). For example, Lagrangian covariance functions introduce advection-specific parameters that can explicitly encode the effect of wind propagation on the magnitude and asymmetry of the resulting space-time correlations (cox1988simple; ma2003families).
@@
An alternative paradigm to geostatistical approaches are the so-called dynamic spatio-temporal models (DSTMs) which describe the conditional evolution of a spatial process over time (wikle2010general). In this context, a prevalent example is the integro-difference equation (IDE) framework, which models the dependence between a spatial process at a future time conditioned on the same process at the present time through an integral operator with a spatial redistribution kernel (wikle1999dimension). IDE models have been proposed to predict spatio-temporal cloud movement (Wikle2002), precipitation (xu2005; liu2022statistical), aerosols (calder2011modeling), and more recently, wind power (ye2024ieee). Both geostatistical models and DSTMs offer distinct advantages: geostatistical approaches are typically regarded as capable surrogate models for complex processes, whereas DSTMs can be highly effective for extrapolating dynamic processes by virtue of their explicit conditional evolution mechanism.
@@
Different in principle than the first two statistical approaches (i.e., geostatistical models and DSTMs), deep learning methods have demonstrated significant promise in wind forecasting applications by developing architectures that are tailored to extract complex patterns from spatio-temporal wind data (dlreview4wind). These include models based on convolutional neural networks (CNNs) (zhu2019learning), recurrent neural networks (RNNs) (ghaderi2017), graph neural networks (GNNs) (GNNwind1; GNNwind2), and more recently, transformer models (bentsen2023spatio). Despite their demonstrated predictive performance, deep learning methods still fall short in terms of uncertainty quantification and interpretability. Thus, recent lines of research seek hybrid methods, referred to collectively as statistical deep learning approaches, in which elements or characteristics of a deep learning architecture are embedded within a statistical model to perform certain learning tasks. For example, deep Kalman filters (krishnan2015deep) and deep state-space models (deepssm) learn the parameters of a state-space model using multi-layer perceptrons (MLPs), and recurrent neural networks (RNNs), respectively. Deep autoregressive models (DeepAR) entail a sequence of RNNs that are structured in an autoregressive-inspired manner (deepar), wheras deep IDE models invoke a CNN to learn the IDE redistribution kernel parameters using past process observations (de2019deep; zammit2020deep).
@@
Departing from the existing literature, we propose a multi-output IDE model characterized by a multivariate, nonstationary redistribution kernel with functional parameters, which is tailored to capture the complex dependencies and interactions of local wind fields across the space-time-height trio. To exploit the wealth of exogenous information available to the forecaster (typically in the form of high-dimensional images), we embed an advanced deep learning architecture within the multi-output IDE model in order to learn key physically meaningful kernel parameters that encode physically relevant information about local wind field dynamics. The proposed model is therefore statistical in nature, enabling full probabilistic inference and forecasting across space, time, and height, with the internalized deep learning architecture acting as a “physics extractor,” to learn a low-dimensional set of physically relevant parameters using high-dimensional data that are difficult to process using standard statistical machinery.
Our proposed model departs from the vast majority of space-time wind forecasting models (be it geostatistical, DSTMs, or deep-learning-based) that adopt a two-dimensional world view of the wind field (space $\times$ time) by focusing on a single representative height. It belongs to the class of statistical deep learning models reviewed above, and especially builds on the deep IDE models therein (de2019deep; zammit2020deep), but differs in two key aspects. First, it generalizes the single-output IDE approaches to the multivariate setting to model the complex and time-varying dependencies across space, time, and height. In that way, the proposed model is regarded as a multi-output IDE, akin to how multi-output GPs generalize their single-output counterparts. The need for the proposed modeling framework is motivated by our data analysis, presented in Section 3, where the multivariate time series representing the multi-height offshore wind speeds are shown to exhibit complex, time-varying cross-correlations, calling for a multi-output dynamical approach in which the height dimension is modeled as a physically meaningful vertical axis rather than merely another spatial coordinate. This multivariate time series approach enables us to learn the complex cross-height dependencies and interactions that cannot be handled by simply treating height as a geospatial coordinate. Another relevant work to this discussion is the multivariate spatio-temporal mixed effects model (MSTM) introduced by mstm, which is designed to model correlated multivariate areal data over space and time. While the MSTM falls under the broader class of multivariate DSTMs, it is not based on the IDE framework, and hence is not designed to model the complex dependencies inherent in the full-fledged wind field across space, time, and height.
@@
Second, unlike existing deep IDE models which use information that is endogenous to the process (e.g., past observations) to train simple deep learning architectures (typically a CNN), our model exploits the wealth of exogenous information available to the forecaster, typically in the form of high-dimensional image streams. For example, when forecasting wind speeds, one has access to large amounts of high-dimensional data from regional numerical weather models about the meso-scale wind velocities, pressure and temperature systems, etc. that encode rich physical information about the local wind dynamics. To fully unlock access to this exogenous information, we propose to embed a transformer model—an advanced deep learning architecture with evidenced success in computer vision and high-dimensional data mining (transformer)—within the proposed multi-output IDE framework as a “physics extractor,” i.e., to learn and update a low-dimensional set of time- and height-varying parameters encoding the physics of local wind field formation and propagation. This not only enables tapping into exogenous, high-dimensional data streams for improved forecasting (instead of being confined to endogenous, historical time series as in previous deep IDE models), but further elevates the role of the embedded deep learning architecture from an opaque black-box to a “physics extractor” by learning key physically meaningful parameters that directly inform the statistically derived forecast.
@@ 3 Data description and analysis
This work is motivated by the potential large-scale offshore wind energy developments in the Northeastern United States, and in particular, the NY/NJ Bight (leasenj). The future offshore wind energy areas in this region, as well as in other locations world-wide, are expected to install ultra-scale offshore wind turbines (hub heights >140>140m and rotor diameters >200>200m) (GE). We make use of two sources of data with varying spatial and temporal resolutions: (i) A set of wind speed measurements collected, at multiple heights, by three floating lidar buoys; and (ii) a set of numerical weather predictions (NWPs), of multiple meteorological variables, obtained from a regional numerical weather model. Details of both sets of data are described next.
Figure 1: (a) Locations of the three sites, E05N, E06, and ASOW6, on top of the offshore wind energy areas in the NY/NJ Bight. The background map is generated using the Northeast Ocean Data Portal (map); (b) Wind rose plot for the wind speeds, averaged over space and height; (c) The empirical temporal covariogram of wind speeds at multiple heights (different colors) and locations (different markers); (d) Time series of wind speeds at all locations (different colors) and heights (different markers), showing strong spatial, temporal, and vertical dependencies.